OpenStudy (anonymous):

What is the product? (5r + 2)(3r − 4) 15r2 + 8 15r2 − 8 15r2 + 14r − 8 15r2 − 14r − 8 Thank you in advance :)

OpenStudy (texaschic101):

you familiar with the FOIL method ?

OpenStudy (dayakar):

(5r + 2)(3r − 4) =5r(3r-4)+2(3r-4)

OpenStudy (anonymous):

no im not @texaschic101

OpenStudy (dayakar):

\[=15r^{2}-20r+6r-8\]

OpenStudy (dayakar):

simplify it

OpenStudy (texaschic101):

FOIL.....first, outside, inside, last (5r + 2)(3r - 4) First -- multiply first terms in each set....5r * 3r = 15r^2 Outside -- multiply outside terms in each set...5r * -4 = -20r Inside -- multiply inside terms in each set...2 * 3r = 6r Last -- multiply last terms in each set...2 * -4 = -8 now put it all together... 15r^2 - 20r + 6r - 8 --- simplify by combining like terms 15r^2 - 14r - 8 <-- your answer any questions ?

OpenStudy (anonymous):

no questions :) thanks to both of you

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

Ill let you know if i need help with something else :)